the width of a rectangle is one half its length. The perimeter of the rectangle is 54 cm. What are the width and length of the rectangle?
Answer by checkley77(12569)

W=L/2
2L+2W=54
2L+2*L/2=54
2L+L=54
3L=54
L=54/3
L=18 ANS. FOR THE LENGTH.
W=18/2=9 ANS. FOR THE LENGTH.
PROOF:
2*9+2*18=54
18+36=54
54=54